OpenStudy (anonymous):

What is the sum of the first 20 terms of the series made from the sequence 6, 9, 12? What is the sum of the first 30 terms of the series made from the sequence in 0.3, 0.1, 0?

OpenStudy (anonymous):

can you identify the pattern? identify the pattern and then write down all the sequence then add them for each question.

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

do you understand what a pattern is or what it means?

OpenStudy (helder_edwin):

u have \[ \large a_1=6 \] \[ \large a_2=9=6+3=a_1+3\cdot1 \] \[ \large a_3=12=9+3=a_2+3=(a_1+3)+3=a_1+3\cdot2 \]

OpenStudy (helder_edwin):

what do u think \[ \large a_{20}= \] is going to be?

OpenStudy (anonymous):

@helder_edwin